OpenStudy (anonymous):

Use continuity to evaluate the limit. (Round your answer to three decimal places.) lim x→3 arctan x^2 − 9 / 3x^2 − 9x

OpenStudy (anonymous):

hint: x^2-9=(x+3)(x-3) and 3x^2 − 9x=3x(x-3) later use definition

OpenStudy (anonymous):

9x on the bottom

OpenStudy (anonymous):

right

OpenStudy (anonymous):

i mean 3x, not 9x